OpenStudy (anonymous):

Find the centroid of the region bounded by the graphs of f(x)= 4-x^2 and g(x)= x+2 Please Help!

OpenStudy (tkhunny):

You'll need some limits. What region is enclosed? Provide both x limits and y limits. You'll need both.

OpenStudy (anonymous):

This problem has no given limits...It just states what I stated above.

OpenStudy (anonymous):

It says to find the centroid of the region bounded by the graphs f(x)= 4-x^2 and g(x)= x+2 Thats all it gives.

OpenStudy (anonymous):

Nothing more...Nothing less.

OpenStudy (anonymous):

to find the centroid u must find the region

OpenStudy (anonymous):

Yes, so how do we do that?

OpenStudy (tkhunny):

You're not paying attention. Do you want them to set up the integrals for you, too? You must use your functions and FIND the limits.

OpenStudy (tkhunny):

They intersect where x and y are equal.